OpenStudy (anonymous):

a pyramid has a height of 16cm and a base with area of 84 sq. cm. what is its volume?

OpenStudy (anonymous):

v of pyramid =1/3(base area)*(hieght) therefore ,,,,=1/3*16*84 =16*28 448 your ans